OpenStudy (anonymous):

is 7/6 a a improper fraction

OpenStudy (displayerror):

A mixed fraction is of the form \[A \frac{B}{C}\] Where B < C. To convert to an improper fraction, you would perform the following operation: \[\frac{A \times C + B}{C}\]